### v2

# Exploring Cross-model Neuronal Correlations in the Context of Predicting Model Performance and Generalizability

Abstract
As Artificial Intelligence (AI) models are increasingly integrated into critical systems, the need for a robust framework to establish the trustworthiness of AI is increasingly paramount. While collaborative efforts have established conceptual foundations for such a framework, there remains a significant gap in developing concrete, technically robust methods for assessing AI model quality and performance. A critical drawback in the traditional methods for assessing the validity and generalizability of models is their dependence on internal developer datasets, rendering it challenging to independently assess and verify their performance claims. This paper introduces a novel approach for assessing a newly trained model’s performance based on another known model by calculating correlation between neural networks. The proposed method evaluates correlations by determining if, for each neuron in one network, there exists a neuron in the other network that produces similar output. This approach has implications for memory efficiency, allowing for the use of smaller networks when high correlation exists between networks of different sizes. Additionally, the method provides insights into robustness, suggesting that if two highly correlated networks are compared and one demonstrates robustness when operating in production environments, the other is likely to exhibit similar robustness. This contribution advances the technical toolkit for responsible AI, supporting more comprehensive and nuanced evaluations of AI models to ensure their safe and effective deployment. Code is available at https://github.com/aheldis/Cross-model-correlation.git.
Introduction
In recent years, artificial intelligence (AI) has emerged as a transformative force, revolutionizing diverse sectors with its unparalleled capabilities. No longer confined to ancillary services, AI applications are increasingly being found in centralized roles of what might be deemed ’essential services’ (Guilherme 2019). The integration of AI models into domains like healthcare, law enforcement, public works, and more, has brought about unprecedented opportunities for innovation, efficiency, personalization, and safety, but in-turn newfound urgency to the challenge of AI safety.
As AI capabilities continue to mature and models are incorporated into increasingly critical aspects of daily-life, it becomes imperative to establish robust techniques for the validation of these models (Batarseh, Freeman, and Huang 2021). This necessitates the development of a robust validation processes to ensure that AI models not only meet but exceed the stringent standards required for deployment in real-world scenarios (Sujan et al. 2023).
This technical challenge goes hand-in-hand with the growing realization from regulatory and licensing agencies that increasing scrutiny over AI is needed (Clark and Hadfield 2019), with a growing chorus of policy analysts advocating for a similar level of rigor applied to the review and approval of AI models that are currently applied to domains like hardware, pharmaceuticals, and medical devices that undergo extensive, standardized, safety reviews prior to their release into public use (O’Sullivan et al. 2019).
Yet, despite the growing criticality of validating AI models, the task presents a formidable challenge. The current state of the art in AI model validation is characterized by complexities and uncertainties, stemming from the inherent nature of the training and validation of machine learning algorithms. A variety of efforts including trials, simulations, model-centered validation, and expert opinion, are employed during training to attempt to validate a model’s performance (Myllyaho et al. 2021). However, what is notable about all these approaches is their dependency on internal developmental factors including a complete reliance on developer-generated training data and agents, to test and validate models for release. Post-market validation is also often present in best-practices, with various methods including failure monitors, safety channels, redundancy, voting, and input and output restrictions being employed to continuously validate the systems after deployment (Myllyaho et al. 2021), but these methods, while crucial, have the drawback of being able to gather data only once a model is already in use, and failures have real-world consequences. More generally, in spite of these considerable efforts, the intricate interplay between data, algorithms, and the dynamic nature of the environments these models are deployed to, poses unique obstacles in ensuring the reliability and generalizability of AI models.
What is ultimately needed for a fully robust validation framework, is the addition of fully external and independent means of evaluating models, an analysis that can be effected independently by wholly unaffiliated researchers, laboring without the need for insights and resources from the model’s originators.
In this paper, we delve into the intricacies of validating AI models, examining the challenges posed by an over-reliance on developer curated data sources and the lack of effective external mechanisms to validate models. We then introduce a novel mechanism for evaluating the performance of a model by contrasting the architectures of models. This paper aims to contribute to the discourse on the critical importance of AI models in society and the pressing need for advancing techniques to validate their efficacy, reliability, and ethical implications in the pursuit of enhancing the safety.
Background
Trustworthy AI
There are multiple definitions for AI. A widely accepted definition originates from IEEE-USA, who defines AI as follows: “Artificial Intelligence (AI) is the theory and development of computer systems that are able to perform tasks that normally require human intelligence such as visual perception, speech recognition, learning, decision-making, and natural language processing” (IEEE USA 2017).
It is particularly in the domains for which AI is called upon for decision-making, that trustworthiness of the models empowered to do so become of paramount importance. There has been considerable work done to-date to specify a precise definition for Trustworthy AI. While multiple definitions exist, for the purposes of this paper we will reference the National Institute of Standards and Technology (NIST) Artificial Intelligence Risk Management Framework, which defined the ”characteristics of trustworthy AI systems include: valid and reliable, safe, secure and resilient, accountable and transparent, explainable and interpretable, privacy-enhanced, and fair with harmful bias managed” (NIST 2023). From this definition we can extrapolate a working definition that Trustworthy AI refers to the development and deployment of artificial intelligence systems that exhibit reliability, transparency, fairness, and ethical considerations throughout their lifecycle.
While all the aforementioned elements of trustworthiness are critical, for the purposes of this paper, we will focus on the Accuracy/performance and robustness elements of a model’s trustworthiness.
Accuracy and Performance
The primary goal of any AI model is to accurately and effectively perform the tasks for which it was designed. This involves achieving high accuracy on relevant metrics, minimizing errors, and demonstrating consistent performance across diverse datasets. NIST speaks to this, formally defining accuracy as “closeness of results of observations, computations, or estimates to the true values or the values accepted as being true” (NIST 2023).
Robustness and Generalization
An AI model’s validity extends beyond its performance on the training data to its ability to generalize well to new, unseen data. Robust models maintain accuracy even in the presence of noise, outliers, or variations in the input data distribution. NIST defines robustness or generalizability more concisely as the “ability of a system to maintain its level of performance under a variety of circumstances” (NIST 2023).
Robustness requires not only that the system perform exactly as it does under expected uses, but also that it should perform in ways that minimize potential harms to people if it is operating in an unexpected setting.
Validity and reliability for deployed AI systems are often assessed by ongoing testing or monitoring that confirms a system is performing as intended. Measurement of validity, accuracy, robustness, and reliability contribute to trustworthiness and should take into consideration that certain types of failures can cause greater harm.
Model Validity
Despite the recognized importance of establishing the validity of AI models in generlized contexts (i.e., that a model will continue to perform to a predetermined level of quality outside it’s training domain), defining an objective, universally applicable measure is a challenging task.
Even setting as an assumption that different stakeholders who may prioritize various aspects of model performance differently can settle upon a universally accepted set of validation metrics that encapsulate all relevant dimensions of validity, the validity of an AI model is also context-dependent, varying based on the application domain and the specific requirements of users. A one-size-fits-all approach to validation may simply not be feasible, given the diverse applications of AI across sectors (Panagoulias, Virvou, and Tsihrintzis 2022).
More intrinsically, a critical limitation when determining the universal applicability of a model is its general reliance on the dataset it was trained on (Tao, Gao, and Wang 2019). These data, often collected and curated by the model-developer themselves, is generally comprised of a highly specific set of features and transformations, collected under unique circumstances, that are virtually impossible in many instances to recreate independently. This makes it difficult to replicate and reproduce the model independently, a hallmark of the peer-review process in most disciplines.
Furthermore, owing to Machine Learning’s overriding goal of maximizing predictive performance, and not establishing any statistically rigorous assertion of an underlying relationship between features and labels, we do not typically have the capacity to leverage statistical hypothesis testing to assert the generalizability of our models. Validation efforts can try to some statistical rigor through simulations (Myllyaho et al. 2021) but such simulations remain dependent on the initial dataset to define the distribution of feature-values in an augmented dataset.
To be able to assert that performance observed on a (typically) small subset of training data will continue to bear true outside of the training environment, researchers typically default to the classic Train/Test split where data is purposefully withheld from the model training in order to test it’s robustness against data it hasn’t previously encountered as a mechanism to infer its generalizability (Tao, Gao, and Wang 2019). While this is often an effective proxy, and tools like cross-validation, or model manipulations like regularization, are employed to further help reinforce this approach, its underlying validation’s dependence on the dataset it was trained on, cannot be fully eliminated, nor can the broader effort to achieve independent verifiability or reproducibility be achieved.
In summary, while there is some consensus on what the critical elements of a robust AI model validation might be, the inherent subjectivity, context dependence, and dependence on limited datasets, pose formidable challenges in implementing an objective and universally applicable measure of validity.
One potential approach to addressing this challenge could emerge, not from the data used to train the model, but rather the resulting model architecture itself. While the formation of a model architecture is almost entirely data dependent, we have some inclination that the structure of the model itself (e.g., an overtrained model with high variance), could yield valuable insights into the performance of that model.
Taken one step further, we explore the potential of studying one model, a relatively unknown one, against the context of another, known and validated model to ascertain whether similarities in neuronal-level correlation, can be sufficient to determine aspects of robustness.
Model Robustness
Neural networks have been widely applied in security applications including spam and phishing detection, intrusion prevention, and malware detection. This black-box method, however, where models are deployed without an understanding for their underlying mechanisms of discriminant classification, is shown to be vulnerable to adversarial attacks (Szegedy et al. 2014). There are two types of adversarial attacks: white-box and black-box. In white-box attacks, the attacker has complete knowledge of the target model and data. Among the existing white-box attack techniques, Fast Gradient Sign Method (FGSM) (Liu, Yu, and Su 2019), DeepFool (Moosavi-Dezfooli, Fawzi, and Frossard 2016), and Projected Gradient Descent (PGD) (Madry et al. 2018) are the most prominent and widely adopted methods. These techniques primarily leverage the gradient of the loss function in their algorithm (Costa et al. 2024). Conversely, in black-box attacks, only the model weights and data are accessible. Hence, some black-box methods opt to utilize features extracted from images instead of relying on model properties (Wicker, Huang, and Kwiatkowska 2018).
Many of the proposed adversarial attacks including per-sample and universal have demonstrated the ability to transfer and generalize effectively to unknown models targeting either classification or interpretation (Costa et al. 2024; Oskouie and Farnia 2022). This fact is notable as these models are trained on completely independent data, and often focus on entirely unique domains. That this transferability is possible is largely attributed to the inherent underlying structural similarities among various neural networks, which arise from the common use of gradient descent-based optimization techniques during the training process (Costa et al. 2024).
The apparent success of cross-model attack techniques implies that the potential exists to establish a correlation score between neural networks, and more specifically, between well-established neural networks with more extensively audited and validated performance, and newly trained models lacking a similarly robust track record. Were such a correlation established, it could provide us with meaningful insights into the potential robustness and generalizability of a newly trained neural network prior to taking it into production.
To that end, in this paper, we introduce a novel methodology to accurately estimate this metric. We detail the technical approach for establishing a cross-model neuronal correlation. Next, we explore the implications of this score in validating an unknown model by leveraging the validation of another established model.
Methods
The correlation between the output vectors of two neurons $N_{i}$ and $N_{j}$ with respect to the input vector $X$, is determined by the Pearson correlation coefficient $\rho(.,.)$ computed by
$\rho(A,B)=\frac{cov(A,B)}{\sigma_{A}\sigma_{B}},$ (1)
where $cov(.,.)$ denotes the covariance function and $\sigma_{.}$ is the standard deviation.
As suggested by (Jin, Yi, and Huang 2022), the neuronal correlation of the $l$-th layer’s representation $R_{l}$ is
$Corr(R_{l})=\frac{1}{n(n-1)}\sum_{i,j=1;\ i\neq j}^{n}|\rho(R_{li},R_{lj})|,$ (2)
where $i$ shows the activation of the $i$-th neuron and $n$ is the dimension of $R_{l}$.
This existing methodology is limited to comparing correlations within a neural network and cannot provide a representation of two distinct neural networks. To extend the definition to allow for correlations across networks, we iterate through each neuron $N_{i}$ in the first network and determine the most correlated neuron $N_{j}$ in the second network. The correlation score of $N_{i}$ denoted as $S_{i}$ is derived by
$S_{i}=\frac{\rho(N_{i},N_{j})}{|\text{layer}(N_{i})-\text{layer}(N_{j})|+1}.$ (3)
The inclusion of the denominator in the formula is crucial, as it accounts for the expected decrease in correlation between neurons that are more distant from each other in the respective architectures. The score is estimated for all neuron pairs. Following that, the average over all $S_{i}$ and $S_{j}$ values determines the correlation score between the two neural networks in question. Given that $\rho$ falls within the interval of -1 to 1, the resultant correlation score similarly spans from -1 to 1. It should be noted that we calculate the correlation function specifically for neurons in the hidden layer, as the last linear layer does not contribute significantly to understanding the network’s behavior and interpretability.
Computing the correlation between all neurons in two networks can be extremely time-intensive, especially for larger architectures. To address this challenge, we also propose a more efficient approach which involves calculating partial correlations, particularly for networks that share some structural similarities. For instance, when comparing two networks with differing layer counts but matching output sizes in certain layers, the correlation analysis can be confined to these comparable layers. While this method sacrifices some accuracy, it offers a significant improvement in computational efficiency, making it more practical for analyzing larger and more complex neural networks.
Additionally, we explore the vulnerability of neural networks to black-box adversarial attacks and examine how this susceptibility relates to our proposed metric. To investigate the architectural relationships between different neural networks, we adapted the attack method introduced by (Papernot et al. 2017), with slight modifications to their original approach. This black-box method facilitates model comparison without necessitating prior knowledge of their architectures or training datasets. As described in Algorithm 1, at each iteration, we train a simple model like a two-layer CNN that differs from the other models used in the experiments on the dataset. Subsequently, we adjust the data using the perturbation computed via the Fast Gradient Sign Method (FGSM) (Liu, Yu, and Su 2019). We iterate through the process multiple times to generate our ultimate adversaries. In our method, we solely utilize test data instead of custom-crafted data, as the former remains untouched during training. Furthermore, we refrain from altering labels based on the oracle’s feedback and discard intermediate adversaries generated throughout the process, retaining only the final iteration to reduce the time and space complexity.
Algorithm 1 Substitute DNN Training
  Input: Training data $D=(x_{i},y_{i})_{i=1}^{n}$, simple two-layer CNN $C$, perturbation norm $\epsilon$
  Initialize perturbed data $P_{0}$ with $X$
  for epoch $\in\ \{1,\dots,N_{ep}\}$ do
     Train $C$ on $P_{epoch-1}$
     Update $P_{epoch}\leftarrow\{\bigl{(}\tilde{x}+\epsilon\cdot\text{sgn}(J_{C}(\tilde{x}),y\bigr{)}:(\tilde{x},y)\in P_{epoch-1}\}$
  end for
  Return $P_{N_{ep}}$
Numerical Results
[…]
Figure 1: Test images and their predicted classes before and after attack.
Table 1: Description of FCNs.
Network Number of neurons in hidden layer
Net1 [100]
Net2 [60, 40]
Net3 [40, 60]
Net4 [196, 392]
Net5 [20, 20, 20, 20, 20]
Table 2: Correlation between FCNs.
(a) MNIST
Net1 Net2 Net3 Net4 Net5
Net1 1 0.644 0.646 0.642 0.534
Net2 X 1 0.697 0.699 0.564
Net3 X X 1 0.730 0.551
Net4 X X X 1 0.541
Net5 X X X X 1
(b) Fashion-MNIST
Net1 Net2 Net3 Net4 Net5
Net1 1 0.684 0.666 0.671 0.614
Net2 X 1 0.724 0.724 0.664
Net3 X X 1 0.747 0.639
Net4 X X X 1 0.604
Net5 X X X X 1
(c) Kuzushiji-MNIST
Net1 Net2 Net3 Net4 Net5
Net1 1 0.572 0.556 0.581 0.510
Net2 X 1 0.623 0.619 0.512
Net3 X X 1 0.679 0.538
Net4 X X X 1 0.530
Net5 X X X X 1
(d) Afro-MNIST
Net1 Net2 Net3 Net4 Net5
Net1 1 0.576 0.573 0.588 0.373
Net2 X 1 0.695 0.705 0.500
Net3 X X 1 0.699 0.485
Net4 X X X 1 0.458
Net5 X X X X 1
Our numerical experiments were designed to investigate the performance of our proposed correlation method across two scenarios: small neural networks and large neural networks. This approach allowed us to comprehensively evaluate the method’s efficacy and scalability.
A. Small neural networks
We evaluated our method on five simple fully connected networks (FCNs) and a two-layer convolutional neural network (CNN). The number of neurons considered for each hidden layer is described in Table 1. These networks were trained on four datasets including MNIST (LeCun 1998), Fashion-MNIST (Xiao, Rasul, and Vollgraf 2017), Kuzushiji-MNIST (Clanuwat et al. 2018), and Afro-MNIST (Wu, Yang, and Prabhu 2020). All these datasets contain 60000 training examples and 10000 test examples. Each example comprises a $28\times 28$ image paired with a classification label from a set of 10 categories. MNIST, Fashion-MNIST, Kuzushiji-MNIST, and Afro-MNIST contain images of handwritten numerals, fashion, Hiragana characters, and African numbers respectively. The selection of datasets with similar attributes was deliberate as our focus was on understanding the interplay between networks, irrespective of data size or category.
Table 3: Accuracy of FCNs before and after attack.
Network Metric MNIST Fashion-MNIST Kuzushiji-MNIST Afro-MNIST
Net1 Accuracy before attack 96.2% 86.8% 84.4% 99.2%
Accuracy after attack 73.7% 59.5% 50.4% 45.1%
Relative degradation 23.4% 31.5% 40.3% 54.5%
Net2 Accuracy before attack 96.6% 86.1% 84.4% 99.4%
Accuracy after attack 65.7% 56.9% 47.4% 44.3%
Relative degradation 32.0% 33.9% 43.8% 55.4%
Net3 Accuracy before attack 96.0% 84.2% 83.8% 99.1%
Accuracy after attack 69.4% 53.3% 48.2% 44.5%
Relative degradation 27.7% 36.7% 42.5% 55.1%
Net4 Accuracy before attack 96.1% 86.8% 84.5% 99.2%
Accuracy after attack 67.4% 58.1% 47.2% 45.0%
Relative degradation 29.9% 33.1% 44.1% 54.6%
Net5 Accuracy before attack 92.9% 77.5% 74.3% 98.5%
Accuracy after attack 53.5% 59.5% 39.6% 40.3%
Relative degradation 42.4% 23.2% 46.7% 59.1%
CNN Accuracy before attack 98.9% 89.9% 91.7% 100.0%
Accuracy after attack 70.8% 59.5% 33.9% 27.1%
Relative degradation 28.4% 33.8% 63.0% 72.9%
Table 3 illustrates the robustness of FCNs against the black-box attack detailed in Algorithm 1. Effect of the attack on some test images is illustrated in Figure 1. Results for CNN are also provided, showcasing its inferior robustness, given its role in generating the adversaries. In our specific scenario, we selected a norm value of $\epsilon=\frac{10}{255}$ and $N_{ep}=10$. The results indicate that, on average, Net5, Net2, Net4, Net3, and Net1 ranked from vulnerable to robust. Among the FCNs, Net5 exhibited the lowest overall accuracy. Notably, according to Table 2(d), Net5 displayed the weakest correlation with the other networks, aligning with its subpar performance. Conversely, Net3 demonstrated the highest average correlation score with the other networks, potentially attributed to its architectural similarity with the other FCNs.
In Table 4, the correlation between FCNs and CNN is investigated. Here, due to the large number of neurons involved, we had to limit the correlation analysis to only $10$ test data points. Looking at the correlation values between CNN and the different networks, Net3 consistently demonstrates the highest correlation values across multiple datasets. This high correlation between Net3 and CNN suggests that their predictive patterns and feature extraction align closely across various datasets. In addition, it’s noteworthy that Net5 exhibits the lowest correlation among the networks, which aligns with expectations due to its inferior performance and robustness.
Table 4: Correlation between CNN and other networks.
Dataset Net1 Net2 Net3 Net4 Net5
MNIST 0.880 0.856 0.908 0.831 0.416
Fashion-MNIST 0.756 0.822 0.820 0.755 0.317
Kuzushiji-MNIST 0.900 0.769 0.910 0.894 0.519
Afro-MNIST 0.588 0.629 0.730 0.718 0.059
Table 5: Partial correlation between ResNets.
ResNet-18 ResNet-34 ResNet-50 ResNet-101 ResNet-152
ResNet-18 1 0.639 0.203 -0.263 0.054
ResNet-34 X 1 0.267 -0.349 0.015
ResNet-50 X X 1 -0.534 -0.407
ResNet-101 X X X 1 -0.688
ResNet-152 X X X X 1
B. Large neural networks
The proposed correlation metric was also investigated on different variants of ResNet, including ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 (He et al. 2016). We used the pretrained weights on ImageNet (Deng et al. 2009) as our baselines. Since investigating the correlation on all of the neurons in all layers was not possible we only tested the method on the first output of the fourth layer on all ResNets. Also, we used only 10 test data points from the validation set of ImageNet. Considering the absolute values of the correlations, results suggest that the most similar networks to ResNet-18, ResNet-34, ResNet-50, ResNet-101, and ResNet-152 are ResNet-34, ResNet-18, ResNet-101, ResNet-152, and ResNet-101. This pattern illustrates that networks with a similar number of layers exhibit stronger partial correlations, serving as a validation of the effectiveness of our proposed correlation metric.
Discussion
These experimental results indicate that our proposed correlation score can effectively demonstrate connections and relationships across diverse neural networks. While these findings are still preliminary, we assert that this score can potentially offer insights into various factors including:
  • Architectural size impact: Smaller architectures tend to exhibit higher correlation with other networks due to their simplicity and shared properties with most neural networks. As networks increase in complexity, this correlation diminishes, reflecting the nuanced decision-making processes that diverge from standard neural network behaviors. Calculating partial correlations, especially for the final layers of the networks, can yield more interpretable insights.
  • Performance implications: Lower correlations may correlate with poorer model performance, indicating that neuron activations deviate from the expected patterns necessary for accurate decision-making.
  • Robustness: A strong correlation with a robust model may suggest a level of robustness within another model as well.
We further note that despite the valuable insights provided by the proposed correlation method, a principle drawback lies in the suboptimal performance in terms of time complexity for obtaining this score. Moreover, when dealing with larger models, a more efficient approach for calculating correlation with improved time complexity and precision may be necessary. Another limitation of our current methodology is the inability to pinpoint the exact reasons behind a low correlation score. However, if a high correlation is observed between two networks, where one is recognized for its accuracy and robustness, there is a strong indication that the other network is also performing well and likely to generalize. Nevertheless, this inference is not entirely precise.
Conclusion
In the ongoing effort to establish a robust framework for AI trustworthiness, a critical component of any eventual solution will be analyses that can be performed by objective and independent researchers without the need for the data the model was trained with initially. One possible approach may be to examine the architecture of the model in question in reference to another established model, in the hopes of gleaning insights into robustness and generalizability. We proposed an approach of taking an aggregated correlation score of neuronal activation across two independent neural networks. Our resulting correlation score effectively reveals connections across diverse neural networks. This score highlights performance implications, suggesting that lower correlations may signal deviations from expected activation patterns, and further indicates that strong correlations with robust models imply similar expectations for performance in other networks. Additionally, our method addresses a critical gap in responsible AI development by providing a robust approach to evaluate model quality and generalizability, independent of internal developer datasets. In spite of these promising results, this approach faces limitations in precision with larger models and cannot pinpoint particular reasons for low correlation scores, making inferences less specific. Future research should focus on resolving the deficiencies, optimizing efficiency, and enhancing diagnostic capabilities to deepen insights into neural network behaviors, ultimately supporting the safe and effective deployment of AI systems.